OpenStudy (anonymous):

what is the equation of the line passing through the points (-2,2) and (3,-1) a. y=-5/3x+4/3 b. y=x-4 c. y=-3/5x+4/5 d. y=-3x+8

OpenStudy (anonymous):

do you know the formula

OpenStudy (anonymous):

The general form of a line is y = mx + b. Using this, you can find your answer. It really helps to plot the points on a plane.

OpenStudy (anonymous):

\[\LARGE y-y_1=\frac{y_2-y_1}{x_2-x_1}(x-x_1)\] \[\LARGE y-2=\frac{-1-2}{3-(-2)}(x-(-2))\] \[\LARGE y-2=\frac{-3}{5}(x+2)\] \[\LARGE y-2=\frac{-3}{5}x-\frac{6}{5}\] \[\LARGE y=-\frac{3}{5}x-\frac{6}{5}+2\] \[\LARGE y=-\frac{3}{5}x-\frac{6}{5}+\frac{10}{5}\] \[\LARGE y=-\frac{3}{5}x-\frac{4}{5}\]

OpenStudy (anonymous):

do you know the answer?

OpenStudy (anonymous):

you're kidding right?? :)

OpenStudy (anonymous):

sorry my bad, it's not - it should be + this is final row \[\LARGE y=-\frac35x+\frac45\]

OpenStudy (anonymous):

okay got! :)